OpenStudy (anonymous):

10,000,000,000 = 10^10 10^11 1^11 1^10

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (timaashorty):

Is it written like 10^10, 10^11 Or it's just like that?

OpenStudy (anonymous):

10^10

OpenStudy (timaashorty):

Well do you know What 10^10 equal? 10 x 10 x 10 x 10 x 10 x 10 x 10 x 10 x 10 x 10 =

OpenStudy (skullpatrol):

$$\Huge 10^1=10,$$$$\Huge 10^2=100,$$$$\Huge10^3=1,000$$

OpenStudy (timaashorty):

It's not 1^10 for sure. 1 x 1 x 1 x 1 x 1 x 1 x 1 x 1 x 1 x 1 = the answer will still remain 1

OpenStudy (anonymous):

like those little number skull put its like that

OpenStudy (skullpatrol):

$$\Huge 10^n=1\underbrace{000...0}$$

OpenStudy (timaashorty):

Those numbers are exponents. It means to multiply the number as much times as the number of exponent is, For example \[2^3 = 2\ x \ 2 \ x \ 2 = 8\] \[10^2 = 10 \ x \ 10 = 100\] See what I did? for the first one I multiple the 2 three times. And for the second I multipled the 10 two times

OpenStudy (skullpatrol):

$$\Huge\text{n zeros}$$

OpenStudy (skullpatrol):

Your answer 10^10 is correct :-)

OpenStudy (anonymous):

Aweosme

OpenStudy (timaashorty):

Based on this problem you need to take your answers. \[10^{10} = 10 \times 10 \times 10\times 10\times10\times10 \times10\times 10\times 10 \times 10 = 10,000,000,000 \] \[10^{11} = 10 \times 10 \times 10\times 10\times10\times10 \times10\times 10\times 10 \times 10 \times 10 = 100,000,000,000 \] \[1^ {11} = 1 \times 1 \times 1 \times 1\times 1 \times 1\times 1\times 1 \times 1 \times 1 \times 1 = 1 \] \[1^ {10} = 1 \times 1 \times 1 \times 1\times 1 \times 1\times 1\times 1 \times 1 \times 1 \ = 1 \]

OpenStudy (timaashorty):

Did you see what I did @Semperfi ? I know you wanted the answer. But it's better if you learn it especially since it's pretty simple. (: